OpenStudy (anonymous):

Can somebody help me double check this? You are on the surface of the moon and throw a rock straight up. Height after t seconds will be: h(t) = V0*t -1.625t^2 meters. V0 = initial velocity. Q: When does the rock hit the ground/ surface of the moon? (answer is a value of t dependent on V0) Thanks

OpenStudy (mertsj):

\[V _{o}-1.625t^2=0\] \[-1.625t^2=-V _{_{o}}\] \[t^2=\frac{V _{o}}{1.625}\] \[t=\sqrt{\frac{V _{o}}{1.625}}\]

OpenStudy (anonymous):

huzzah. We match. Thanks.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Actually, this kept bothering me. Should be: t = v0/1.625 ... h(t) = V0*t -1.625t^2 ... 1. set equation equal to zeo 2. Factor out a t ...t(V0 - 1.625t) ... t = v0/1.625

OpenStudy (mertsj):

Oh gees. I didn't see that t. Yes. Just like you have it.